OpenStudy (anonymous):

What is the first step in solving the equation eight d minus six equals two for d? Add 6 Subtract 6 Divide by 8 Multiply by 8

OpenStudy (roadjester):

it depends on if there are parentheses

OpenStudy (anonymous):

what would you guess it is. write the equation down first... 8d - 6 = 2 whats an easy thing you can do straight away?

OpenStudy (roadjester):

order of operations

OpenStudy (anonymous):

Multiply by 8

OpenStudy (roadjester):

incorrect however it depends on if it is 8(d-6) =2 OR 8d-6=2

OpenStudy (roadjester):

there is a difference

OpenStudy (anonymous):

well you could do that, that would get the d on its own, but it would mean introducing two more 8's, one multiplying the 6, one multiplying the 2. if you first add 6 then you can get 8d = 2 + 6 8d = 8 which is way easier than your way which would give 8d = 8 x -6 = 8 x 2

OpenStudy (anonymous):

so you divide

OpenStudy (roadjester):

@JamesWolf can you really make the assumption that there are no parentheses? he spelled it out, he didn't use symbols

OpenStudy (anonymous):

yes i can. no you dont divide. you add 6. try both out, dividing 8, and adding 6, see which one simplifies it

OpenStudy (anonymous):

ok for roadjesters sake is it 8(d-6) = 2 or 8d - 6 = 2 ?

OpenStudy (anonymous):

he i am not a he @roadjester and james thank you for helping me

OpenStudy (anonymous):

@Jameswolf

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

though roadjester was trying to help too

OpenStudy (roadjester):

alright, then i apologize however I was only trying to help

OpenStudy (anonymous):

it looks like your name is michael, not michaela

OpenStudy (anonymous):

thank you to @Roadjester

OpenStudy (roadjester):

you're welcome @michaelaepalmore

OpenStudy (anonymous):

A movie rental program charges ten dollars per month plus an additional three dollars for each movie rented. Which equation could be used to determine how many movies were rented if your monthly bill was twenty-two dollars? three plus ten m equals twenty-two ten plus three m equals twenty-two twenty-two plus ten m equals three twenty-two plus three m equals ten can yall help me with this one to plz

OpenStudy (chrisdbest):

10m + 3r = 22 Thats the equation were working with

OpenStudy (chrisdbest):

22 plus 10m equals three or

OpenStudy (anonymous):

thank you @chrisdbest

OpenStudy (chrisdbest):

medal me for the honor

OpenStudy (anonymous):

i cant medal you sorry